Coves Cakes is a local bakery. Price and cost information follows: $ 14.11 Price per cake Variable cost per cake IngredientsRequired 1 Required 2 Required 3 Determine Coves break-even point in units and sales dollars. (Round your Break-Even Units aRequired 1 Required 2 Required 3 Determine the bakerys margin of safety if it currently sells 500 cakes per month. (Round yoRequired 1 Required 2 Required 3 Determine the number of cakes that Cove must sell to generate $2,700 in profit. (Round your
Answer #1

Correct Answer:

Break Even in Volume (units)

                                            460.00

Breakeven sales in dollars

$                                     6,490.60

margin of safety on selling 500 units

$    564

% of margin of safety

8.00%

Volume (units) required to earn $16000 profit

721

Working;

Calculation of Contribution per unit

A

Selling price per unit

$                                           14.11

B

Variable cost per unit

(2.34 + 1.18 + 0.23 )

$                                              3.75

C=A-B

Contribution margin per unit

$                                           10.36

D

Fixed Cost

$                                     4,765.60

Calculation of Margin of safety

D

Fixed Cost

$                                     4,765.60

C

Contribution margin per unit

$                                           10.36

E =D/C

Break Even in Volume (units)

                                            460.00

F= E * A

Breakeven sales in dollars

$                                     6,490.60

G = 500 * A

Sales value at 500 units

$                                     7,055.00

H=G-F

margin of safety

$                                               564

I =H/G *100

% of margin of safety

8.00%

J

Required profit

$                                     2,700.00

D

Fixed Cost

$                                     4,765.60

K=D+J

Contribution margin required

$                                     7,465.60

C

Contribution margin per unit

$                                           10.36

E=K/C

Volume (units) required to earn $ 2700

profit

720.62

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks.